Suppose you are given a 95% confidence interval estimate of monthly income for staff at Texas State. The lower bound of the estimate is $4,950 and upper bound is $7,560. What is the margin of error that was used to compute the interval?

  • A. $1,305

  • B. 0.4750

  • C. $2,610

  • D. $6,255

  • E. 0.9500
Answer #1

x= Sample income estimated monthly X ME x + ME 4950 7560 where ME = margin of enor 2 gives ( x +ME)=4950 - 7560 X-ME) 2 ME 2